# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ Abstract
Autonomous AI agents powered by large language models (LLMs) with structured function-calling interfaces have dramatically expanded capabilities for real-time data retrieval, complex computation, and multi-step orchestration. Yet, the explosive proliferation of plugins, connectors, and inter-agent protocols has outpaced discovery mechanisms and security practices, resulting in brittle integrations vulnerable to diverse threats. In this survey, we introduce the first unified, end-to-end threat model for LLM-agent ecosystems, spanning host-to-tool and agent-to-agent communications, formalize adversary capabilities and attacker objectives, and catalog over thirty attack techniques. Specifically, we organized the threat model into four domains: Input Manipulation (e.g., prompt injections, long-context hijacks, multimodal adversarial inputs), Model Compromise (e.g., prompt- and parameter-level backdoors, composite and encrypted multi-backdoors, poisoning strategies), System and Privacy Attacks (e.g., speculative side-channels, membership inference, retrieval poisoning, social-engineering simulations), and Protocol Vulnerabilities (e.g., exploits in Model Context Protocol (MCP), Agent Communication Protocol (ACP), Agent Network Protocol (ANP), and Agent-to-Agent (A2A) protocol). For each category, we review representative scenarios, assess real-world feasibility, and evaluate existing defenses. Building on our threat taxonomy, we identify key open challenges and future research directions, such as securing MCP deployments through dynamic trust management and cryptographic provenance tracking; designing and hardening Agentic Web Interfaces; and achieving resilience in multi-agent and federated environments. Our work provides a comprehensive reference to guide the design of robust defense mechanisms and establish best practices for resilient LLM-agent workflows.
@@ I Introduction
Autonomous AI agents have advanced rapidly in the past two years, driven by large language models (LLMs) that can perceive, reason, and act in complex environments. The introduction of structured function calling in 2023 enabled LLMs to invoke external APIs in a consistent, machine-readable format, allowing them to fetch real-time data, perform computations, and orchestrate multi-step workflows [1, 2]. At the same time, a vibrant ecosystem of plugins and extension frameworks emerged. ChatGPT plugins provided turnkey access to third-party services; frameworks such as LangChain and LlamaIndex offered standardized connector interfaces; and dedicated AI app stores began to host curated collections of tools [3]. Together, these developments have greatly expanded the practical capabilities of autonomous agents [4, 5, 6]. Despite this progress, tool integration remains siloed and labor-intensive. Developers must write bespoke adapter code for each new service, manually manage credentials and error handling, and learn platform-specific function-calling conventions. Workflows are often hard-coded, which limits an agent’s ability to adapt at runtime and makes extension cumbersome. Moreover, the absence of a shared discovery mechanism or common vocabulary for tool capabilities increases the risk of security gaps when onboarding new services or updating existing ones [7, 8, 9, 10].
@@
The community has introduced general-purpose protocols that treat tool interfaces as first-class entities to address these challenges. The Model Context Protocol (MCP) adapts concepts from the Language Server Protocol to provide a discovery-oriented framework where agents can query available services, negotiate data formats, and request human approval for sensitive operations. Building on MCP, related specifications such as the Agent Network Protocol (ANP) and Agent-to-Agent (A2A) protocol aim to support peer-to-peer collaboration and multi-agent orchestration [11, 12]. Many open-source MCP servers now integrate models with version control systems, cloud databases, and even 3D design suites, illustrating the promise of an extensible and interoperable agent layer [13, 14, 15].
Authors Year Specific Attack ASR (%)
Zhan et al.[16] 2025 Adaptive Indirect Prompt Injection $\geq 50$
Jiang et al.[17] 2025 Compositional Instruction Attack >95>95
Yu et al.[18] 2023 Jailbreak Fuzzing >90>90
Schwartz et al.[19] 2025 Graph of Attacks with Pruning $92$
@@
Chen et al.[20] 2025 Active Environment Injection Attack $93$
Huang et al.[21] 2023 Composite Backdoor Attack $100$
@@
Zhu et al.[22] 2025 Encrypted Multi-Backdoor Implantation $\approx 100$
Zou et al.[23] 2024 Retrieval Poisoning $90$
@@
TABLE I: Attack Success Rates (ASR) reported across selected studies
@@
The alarmingly high success rates summarized in Table I illustrate just how vulnerable LLM‐powered agent workflows have become to input manipulation and protocol exploits. Adaptive prompt injection attacks can bypass existing defenses in over 50% of cases [16], while sophisticated jailbreak techniques routinely achieve greater than 90% ASR—even reaching near‐perfect success for backdoor implants such as CBA and DemonAgent [21, 22]. More strikingly, fuzzing‐based jailbreaks like GPTFuzz and pruning‐based strategies (GAP) report ASRs above 90% [18, 19], and environment‐injection attacks against mobile‐OS agents reach up to 93% [20]. These findings—from prompt injections to deep protocol‐level attacks—underscore the pressing need for holistic threat models and robust defense mechanisms in LLM‐agent workflows, which is the focus of this paper.
@@
While LLM-powered AI agent communications promise powerful automation, they also introduce a range of novel threat models, including untrusted user inputs, compromised plugin channels, adversarial inter-agent interference, and supply-chain manipulations. Research in this area has been fragmented, with individual studies examining isolated exploits without a consolidated framework to understand how these threats interrelate across the full communication stack [24, 25, 26]. Fig. 2 presents the proposed high-level taxonomy of threat models for LLM-powered agents, grouping adversarial strategies into four main categories: Input Manipulation, Model Compromise, System & Privacy, and Protocol Vulnerabilities. This survey fills that gap by first articulating comprehensive threat models for AI agent ecosystems, defining adversary capabilities, attacker goals, and exploitable interfaces at each layer of the protocol and tooling landscape. By systematically cataloging and classifying attack techniques under these threat models, we provide researchers and practitioners with a unified reference that highlights critical vulnerabilities and informs the design of more resilient defense strategies. The main contributions of this study are:
@@
  • We provide a comprehensive taxonomy of input manipulation attacks, covering direct and compositional prompt injections, adaptive hijacks, long-context jailbreaks and multimodal adversarial perturbations.
@@
  • We present an in-depth analysis of model compromise attacks, including prompt-level backdoors, model-parameter backdoors, composite and encrypted multi-backdoors, as well as data and memory poisoning techniques.
@@
  • We review system and privacy attacks that target federated and multi-agent deployments, such as speculative side-channels, membership inference, retrieval poisoning, and social-engineering simulations, illustrating how adversaries can exfiltrate or corrupt sensitive data.
@@
  • We analyze the emerging threat landscape in protocol-based agent communications, focusing on vulnerabilities in both the host-tool and agent-agent layers of MCP and A2A interactions.
@@
  • We identify open challenges and future research directions, highlighting where current defenses fall short from MCP protocol vulnerabilities and agentic web interfaces to memory-centric LLM risks and federated/multi-agent resilience—and propose promising avenues for securing LLM-powered agent ecosystems.
@@
Figure 1 depicts the overall organization of this survey. In Section II, we review the most relevant prior work. Section III surveys input-manipulation techniques, including prompt injections, code tampering, and multimodal adversarial perturbations, that compromise LLM behavior before it reaches the core model. Section IV analyzes backdoor and poisoning attacks that stealthily embed malicious behaviors within a model’s parameters. Section V explores system-level and privacy attacks targeting communications in LLM-powered AI‐agent frameworks. Section VI surveys the protocol vulnerabilities, including, Model Control-Protocol (MCP), Agent Communication Protocol (ACP), Agent Network Protocol (ANP), and Agent-to-Agent (A2A) Protocol. Section VII outlines open challenges and promising directions for future research. Finally, Section VIII concludes the paper.
@@
Authors Year Focus Area Key Contribution Security Coverage Limitations
@@
Yang et al. [13] 2025 Communication protocols for LLM agents Two-dimensional protocol taxonomy; comparison of context-oriented and inter-agent schemes Security, scalability, and latency comparisons Does not address concrete attack vectors or LLM-specific threats
@@
Hou et al. [25] 2025 Model Context Protocol (MCP) Lifecycle and architecture of MCP; industry adoption Examines protocol-stage security No analysis of protocol abuse or agent-level exploitation
@@
Deng et al. [9] 2025 Security challenges in AI agents Four key security threat categories (execution, input, environment, interaction) Emphasis on conceptual vulnerabilities Does not address LLM-specific communication layers
@@
Wang et al. [27] 2025 Full-stack LLM safety Lifecycle-wide taxonomy covering over 800 papers Model alignment, editing, secure pretraining Lacks specific focus on agent communications or protocols
@@
Yan et al. [5] 2025 LLM-powered multi-agent systems Communication roles, strategies, and challenges in MAS Scalability and coordination threats Limited emphasis on attack surfaces or defenses
@@
Chen et al. [28] 2025 Computer-Using Agents (CUAs) Safety analysis taxonomy for GUI-interacting agents Defense categorization and benchmarking Focused more on GUI-level threats than multi-agent protocols
@@
Ferrag et al. [2] 2025 Benchmarks and frameworks for LLM agents Taxonomy of 60+ benchmarks; toolkits and protocol overview Coverage of ACP, MCP, A2A No systematic analysis of vulnerabilities or attacks
@@
Friha et al. [29] 2024 Edge Intelligence + LLMs Analysis of architectures, optimizations, and use cases Identifies edge-related vulnerabilities Does not cover agent coordination or protocol-specific risks
@@
Ferrag et al. [30] 2025 LLMs in cybersecurity workflows Evaluation of 42 LLMs on cyber tasks; dataset overview Covers prompt injection, poisoning, adversarial use Lacks in-depth exploration of agent-based communication threats
@@
This Survey Vulnerabilities in LLM agent communication Threat taxonomy across layers (MCP, A2A); attacks on CrewAI, LangChain, AutoGen Systematic coverage of 30+ attack vectors with practical implications First to focus on end-to-end vulnerabilities in multi-agent LLM protocols
TABLE II: Comparison of Related Surveys on LLMs, Agents, and Security
@@ II-A Communication Protocols and Frameworks
Yang et al. [13] present a timely and in-depth examination of the current limitations and opportunities related to communication protocols for large language model (LLM) agents. As LLM agents become increasingly integrated into domains such as customer service, healthcare, and data analysis, the lack of standardized communication methods poses a significant barrier to scalability and collaboration. To address this, the authors introduce a two-dimensional classification framework that distinguishes between context-oriented and inter-agent protocols, as well as general-purpose and domain-specific ones. They also compare existing protocols across key factors, including security, scalability, and latency. The review paper concludes by identifying critical directions for future research, emphasizing the importance of adaptability, privacy preservation, layered architectures, and mechanisms for enabling group-based interaction to support collective intelligence in multi-agent systems.
Hou et al. [25] provide a thorough analysis of the Model Context Protocol (MCP), a standardized interface developed to facilitate smooth interaction between AI models and external tools or resources. By promoting interoperability and reducing data silos, MCP aims to streamline the integration of models across diverse systems. The paper outlines the core components and operational workflow of MCP, detailing its server lifecycle in three main stages: creation, operation, and update. Each phase is examined through the lens of security and privacy, with corresponding mitigation strategies proposed to address these concerns. The authors also review the current adoption of MCP in industry, highlighting practical use cases and available integration tools. Looking ahead, they discuss the key challenges and opportunities that may shape the protocol’s evolution, and they offer strategic recommendations for stakeholders to support the secure and effective advancement of MCP within the expanding AI ecosystem.

### v2

@@ Abstract
Autonomous AI agents powered by large language models (LLMs) with structured function-calling interfaces have greatly expanded capabilities for real-time data retrieval, computation, and multi-step orchestration. However, the rapid growth of plugins, connectors, and inter-agent protocols has outpaced security practices, leading to brittle integrations—plugin APIs and protocol adapters that rely on ad-hoc authentication, inconsistent schemas, and weak validation—making them vulnerable to failures and exploitation. This survey introduces a unified end-to-end threat model for LLM-agent ecosystems, spanning host-to-tool and agent-to-agent communications, and catalogs over thirty attack techniques across Input Manipulation, Model Compromise, System and Privacy Attacks, and Protocol Vulnerabilities. For each category, we provide a formal mathematical formulation of the underlying threat model, defining attacker capabilities, objectives, and affected layers to enable systematic analysis. Representative examples include Prompt-to-SQL (P2SQL) injections and the Toxic Agent Flow exploit in GitHub’s MCP server. For each category, we assess feasibility, review defenses, and outline mitigation strategies such as dynamic trust management, cryptographic provenance tracking, and sandboxed agentic interfaces. The framework was validated through expert review and cross-mapping with real-world incidents and public vulnerability repositories (e.g., CVE, NIST NVD) to ensure practical relevance. Compared to prior surveys, this work provides the first integrated taxonomy bridging input-level exploits and protocol-layer vulnerabilities in LLM-agent ecosystems while introducing formal system definitions for each threat class. Ultimately, it offers actionable insights for securing next-generation AI agents through layered defense and continuous verification. Our work provides a comprehensive reference to guide the design of secure and resilient LLM-agent workflows.
@@
CVE Common Vulnerabilities and Exposures
CWE Common Weakness Enumeration
@@ I Introduction
Autonomous AI agents have advanced rapidly in the past two years, driven by large language models (LLMs) that can perceive, reason, and act in complex environments. The introduction of structured function calling in 2023 enabled LLMs to invoke external APIs in a consistent, machine-readable format, allowing them to fetch real-time data, perform computations, and orchestrate multi-step workflows [sapkota2025ai, ferrag2025llm, amiri2024deep]. At the same time, a vibrant ecosystem of plugins and extension frameworks emerged. ChatGPT plugins provided turnkey access to third-party services; frameworks such as LangChain and LlamaIndex offered standardized connector interfaces; and dedicated AI app stores began to host curated collections of tools [ferrag2025reasoning, heidari2023secure]. Together, these developments have greatly expanded the practical capabilities of autonomous agents [sapkota2025vibe, yan2025beyond, yehudai2025survey, heidari2023internet]. Despite this progress, tool integration remains siloed and labor-intensive. Developers must write bespoke adapter code for each new service, manually manage credentials and error handling, and learn platform-specific conventions for function calls. Workflows are often hard-coded, which limits an agent’s ability to adapt at runtime and makes extension cumbersome. Moreover, the absence of a shared discovery mechanism or common vocabulary for tool capabilities increases the risk of security gaps when onboarding new services or updating existing ones [jiang2025large, tran2025multi, deng2025ai, romero2025agentic, yigit2025generative].
@@
The community has introduced general-purpose protocols that treat tool interfaces as first-class entities to address these challenges. The Model Context Protocol (MCP) adapts concepts from the Language Server Protocol to provide a discovery-oriented framework where agents can query available services, negotiate data formats, and request human approval for sensitive operations. Building on MCP, related specifications such as the Agent Network Protocol (ANP) and Agent-to-Agent (A2A) protocol aim to support peer-to-peer collaboration and multi-agent orchestration [narajala2025enterprise, ehtesham2025survey]. Many open-source MCP servers now integrate models with version control systems, cloud databases, and even 3D design suites, illustrating the promise of an extensible and interoperable agent layer [yang2025survey, chen2025standalone, wei2025survey].
Authors Year Specific Attack Attack Type ASR (%)
Zhan et al.[zhan2025adaptive] 2025 Adaptive Indirect Prompt Injection Input Manipulation $\geq 50$
Jiang et al.[jiang2025deceiving] 2025 Compositional Instruction Attack Input Manipulation >95>95
Yu et al.[yu2023gptfuzzer] 2023 Jailbreak Fuzzing Input Manipulation >90>90
Schwartz et al.[schwartz2025graph] 2025 Graph of Attacks with Pruning Input Manipulation $92$
@@
Chen et al.[chen2025aeia] 2025 Active Environment Injection Attack Protocol Exploit / Environment Attack $93$
Huang et al.[huang2023composite] 2023 Composite Backdoor Attack Protocol Exploit / Backdoor $100$
@@
Zhu et al.[zhu2025demonagent] 2025 Encrypted Multi-Backdoor Implantation Protocol Exploit / Backdoor $\approx 100$
Zou et al.[zou2024poisonedrag] 2024 Retrieval Poisoning Protocol Exploit / RAG Poisoning $90$
@@
TABLE I: Attack Success Rates (ASR) and attack types reported across selected studies.
@@
The alarmingly high success rates summarized in Table I illustrate just how vulnerable LLM‐powered agent workflows have become to input manipulation and protocol exploits. Attack Success Rates (ASR) values are reported as stated in the original publications. These results are not directly comparable due to variations in model architectures, datasets, experimental setups, and defense mechanisms. Adaptive prompt injection attacks can bypass existing defenses in over 50% of cases [zhan2025adaptive, heidari2025securing], while sophisticated jailbreak techniques routinely achieve greater than 90% ASR—even reaching near‐perfect success for backdoor implants such as CBA and DemonAgent [huang2023composite, zhu2025demonagent, heidari2024everything]. More strikingly, fuzzing‐based jailbreaks like GPTFuzz and pruning‐based strategies (GAP) report ASRs above 90% [yu2023gptfuzzer, schwartz2025graph], and environment‐injection attacks against mobile‐OS agents reach up to 93% [chen2025aeia]. These findings—from prompt injections to deep protocol‐level attacks—underscore the pressing need for holistic threat models and robust defense mechanisms in LLM‐agent workflows, which is the focus of this paper.
@@
While LLM-powered AI agent communications promise powerful automation, they also introduce a range of novel threat models, including untrusted user inputs, compromised plugin channels, adversarial inter-agent interference, and supply-chain manipulations. Research in this area has been fragmented, with individual studies examining isolated exploits without a consolidated framework to understand how these threats interrelate across the full communication stack [zhang2025100, hou2025model, yang2025privacy, asadi2025new]. Fig. 2 presents the proposed high-level taxonomy of threat models for LLM-powered agents, grouping adversarial strategies into four main categories: Input Manipulation, Model Compromise, System & Privacy, and Protocol Vulnerabilities. This survey fills that gap by first articulating comprehensive threat models for AI agent ecosystems, defining adversary capabilities and attacker goals, and identifying exploitable interfaces at each layer of the protocol and tooling landscape. Although prior surveys have explored individual dimensions—such as communication protocols [yang2025survey, hou2025model], general AI-agent security [deng2025ai], and lifecycle safety frameworks for LLM [wang2025comprehensive]—these efforts remain fragmented. To the best of our knowledge, this work presents the first comprehensive synthesis and unified taxonomy that integrates multi-layer threats—spanning input manipulation, model compromise, system and privacy, and protocol vulnerabilities—across LLM-powered agent ecosystems. The proposed threat model is intended to serve a broad range of stakeholders across the LLM-agent ecosystem. For developers and system architects, it provides a structured framework for identifying, prioritizing, and mitigating vulnerabilities in plugin APIs, connector frameworks, and inter-agent protocols during design and deployment. For security researchers, it establishes a reproducible foundation for benchmarking red-teaming efforts, developing defense mechanisms, and advancing empirical evaluations of agentic security. Furthermore, for policymakers and standardization bodies, the model provides actionable insights to inform the development of regulatory guidelines, compliance standards, and certification processes for secure and trustworthy AI-agent systems. The proposed threat model is intended for developers, security researchers, and policymakers seeking to identify, assess, and mitigate vulnerabilities in LLM-agent ecosystems.
The key contributions of this work are summarized as follows:
@@
  • We provide a comprehensive taxonomy of input manipulation attacks, covering direct and compositional prompt injections, adaptive hijacks, long-context jailbreaks, and multimodal adversarial perturbations. We formally define the Input Manipulation Threat Model to mathematically characterize their behavior and success conditions.
@@
  • We present an in-depth analysis of model-compromise attacks, including prompt-level backdoors, model-parameter backdoors, composite and encrypted multi-backdoors, and data and memory poisoning techniques. We propose a formal definition of the Model Compromise Threat Model to capture parameter- and data-oriented manipulations.
@@
  • We review system and privacy attacks targeting federated and multi-agent deployments, including speculative side channels, membership inference, retrieval poisoning, and social-engineering simulations, and illustrate how adversaries can exfiltrate or corrupt sensitive data. We introduce a formal definition of the System & Privacy Threat Model.
@@
  • We analyze the emerging threat landscape in protocol-based agent communication, focusing on vulnerabilities in both the host-to-tool and agent-to-agent layers of MCP and A2A interactions. We propose the formal definition of the Protocol-Level Threat Model to represent communication-layer exploits mathematically.
@@
  • We identify open challenges and future research directions, highlighting where current defenses fall short — from MCP protocol vulnerabilities and agentic web interfaces to memory-centric LLM risks and federated/multi-agent resilience — and propose promising avenues for securing LLM-powered agent ecosystems.
To construct the proposed taxonomy of over 30 attack techniques, we adopted a systematic, evidence-driven methodology. First, we conducted an extensive literature review of more than 150 peer-reviewed publications and technical reports from reputable sources such as IEEE, ACM, Springer, Elsevier, and arXiv, covering topics including prompt injection, data poisoning, model backdoors, side-channel inference, and protocol exploitation in LLM-powered systems. The search was performed using combinations of keywords such as “LLM security,” “AI agent vulnerabilities,” “prompt injection,” “model compromise,” “retrieval poisoning,” “tool invocation attacks,” “multi-agent communication,” “protocol exploit,” and “Model Context Protocol (MCP) security.” Second, we incorporated real-world incident analyses and proof-of-concept exploits from open repositories, vulnerability databases (e.g., CVE), and GitHub security advisories. Finally, the framework underwent expert validation through structured review by domain specialists and was cross-mapped to real-world incident data to verify coverage and practical applicability. Each identified attack vector was cross-referenced with representative LLM-agent frameworks—including LangChain, AutoGen, and CrewAI—to assess practical feasibility and layer-specific relevance. This combined literature-based and empirical approach ensures that the taxonomy is comprehensive, reproducible, and grounded in both academic rigor and real-world evidence.
@@
To clarify the scope and depth of this work, we explicitly define it as a taxonomy-driven survey framework. The taxonomy provides a structured organization of over thirty attack techniques across four hierarchical domains—Input Manipulation, Model Compromise, System and Privacy Attacks, and Protocol Vulnerabilities—each accompanied by representative examples and corresponding defenses. This structure balances breadth and analytical depth, allowing comprehensive coverage while maintaining detailed insight into each category’s mechanisms and implications. Figure 1 depicts the overall organization of this survey. In Section II, we review the most relevant prior work. Section III surveys input-manipulation techniques, including prompt injections, code tampering, and multimodal adversarial perturbations, that compromise LLM behavior before it reaches the core model. Section IV analyzes backdoor and poisoning attacks that stealthily embed malicious behaviors within a model’s parameters. Section V explores system-level and privacy attacks targeting communications in LLM-powered AI‐agent frameworks. Section VI surveys the protocol vulnerabilities, including the Model Control Protocol (MCP), Agent Communication Protocol (ACP), Agent Network Protocol (ANP), and Agent-to-Agent (A2A) Protocol. Section VII outlines open challenges and promising directions for future research. Finally, Section VIII concludes the paper.
@@
Authors Year Focus Area Key Contribution Security Coverage Threat Model Category Limitations
@@
Yang et al. [yang2025survey] 2025 Communication protocols for LLM agents Two-dimensional protocol taxonomy; comparison of context-oriented and inter-agent schemes Security, scalability, and latency comparisons Protocol Vulnerabilities (PV) Does not address concrete attack vectors or LLM-specific threats
@@
Hou et al. [hou2025model] 2025 Model Context Protocol (MCP) Lifecycle and architecture of MCP; industry adoption Examines protocol-stage security Protocol Vulnerabilities (PV) No analysis of protocol abuse or agent-level exploitation
@@
Deng et al. [deng2025ai] 2025 Security challenges in AI agents Four key security threat categories (execution, input, environment, interaction) Emphasis on conceptual vulnerabilities Input Manipulation (IM) Does not address LLM-specific communication layers
@@
Wang et al. [wang2025comprehensive] 2025 Full-stack LLM safety Lifecycle-wide taxonomy covering over 800 papers Model alignment, editing, secure pretraining System & Privacy (SP) Lacks specific focus on agent communications or protocols
@@
Yan et al. [yan2025beyond] 2025 LLM-powered multi-agent systems Communication roles, strategies, and challenges in MAS Scalability and coordination threats Protocol Vulnerabilities (PV) Limited emphasis on attack surfaces or defenses
@@
Chen et al. [chen2025survey] 2025 Computer-Using Agents (CUAs) Safety analysis taxonomy for GUI-interacting agents Defense categorization and benchmarking Input Manipulation (IM) Focused more on GUI-level threats than multi-agent protocols
@@
Ferrag et al. [ferrag2025llm] 2025 Benchmarks and frameworks for LLM agents Taxonomy of 60+ benchmarks; toolkits and protocol overview Coverage of ACP, MCP, A2A Protocol Vulnerabilities (PV) No systematic analysis of vulnerabilities or attacks
@@
Friha et al. [friha2024llm] 2024 Edge Intelligence + LLMs Analysis of architectures, optimizations, and use cases Identifies edge-related vulnerabilities System & Privacy (SP) Does not cover agent coordination or protocol-specific risks
@@
Ferrag et al. [ferrag2025generative] 2025 LLMs in cybersecurity workflows Evaluation of 42 LLMs on cyber tasks; dataset overview Covers prompt injection, poisoning, adversarial use Input Manipulation (IM) & Model Compromise (MC) Lacks in-depth exploration of agent-based communication threats
@@
This Survey Vulnerabilities in LLM agent communication Threat taxonomy across layers (MCP, A2A); attacks on CrewAI, LangChain, AutoGen Systematic coverage of 30+ attack vectors with practical implications Unified Framework: IM, MC, SP, PV First to integrate all four threat models (Input Manipulation, Model Compromise, System & Privacy, and Protocol Vulnerabilities) into a unified end-to-end taxonomy for LLM-powered agent ecosystems.
TABLE II: Comparison of Related Surveys on LLMs, Agents, and Security, mapped to the four unified threat model categories: Input Manipulation (IM), Model Compromise (MC), System & Privacy (SP), and Protocol Vulnerabilities (PV).
@@ II-A Communication Protocols and Frameworks
Yang et al. [yang2025survey] present a timely and in-depth examination of the current limitations and opportunities related to communication protocols for large language model (LLM) agents. As LLM agents become increasingly integrated into domains such as customer service, healthcare, and data analysis, the lack of standardized communication methods poses a significant barrier to scalability and collaboration. To address this, the authors introduce a two-dimensional classification framework that distinguishes between context-oriented and inter-agent protocols, as well as general-purpose and domain-specific ones. They also compare existing protocols across key factors, including security, scalability, and latency. The review paper concludes by identifying critical directions for future research, emphasizing adaptability, privacy preservation, layered architectures, and mechanisms that enable group-based interaction to support collective intelligence in multi-agent systems.
Hou et al. [hou2025model] provide a thorough analysis of the Model Context Protocol (MCP), a standardized interface developed to facilitate smooth interaction between AI models and external tools or resources. By promoting interoperability and reducing data silos, MCP aims to streamline model integration across diverse systems. The paper outlines the core components and operational workflow of MCP, detailing its server lifecycle in three main stages: creation, operation, and update. Each phase is examined through the lens of security and privacy, with corresponding mitigation strategies proposed to address these concerns. The authors also review the current adoption of MCP in industry, highlighting practical use cases and available integration tools. Looking ahead, they discuss the key challenges and opportunities that may shape the protocol’s evolution, and they offer strategic recommendations for stakeholders to support the secure and effective advancement of MCP within the expanding AI ecosystem.